You are offered the following investment: investing $500 today and receiving $600 in 5 years. What...

  1. You are offered the following investment: investing $500 today and receiving $600 in 5 years.
  1. What is the implied interest rate for this investment?
  1. You can also deposit the money into a bank account that pays 4% annually. Suppose the risks are similar, which investment would you choose?
Answer #1

Answer: A. Implied interest rate that we have to calculate for this period is

Time period : 5 years

Present value = $500

Future value of the investment after 5 years : $600

using the equation FV = PV(1+r)^(no of years)

we have to calculate this r for the implied interest rate

600 = 500 ( 1+r)^(5)

r = (600/500)^(1/5) -1

r = 3.714%

Answer B. We will choose to deposit in the bank account reason being it is offering higher interest rate to us than previous investment which offers less interest rate to us.